# LLM Reading Tea Leaves: Automatically Evaluating Topic Models with Large Language Models

### v1

@@ Abstract
Topic modeling has been a widely used tool for unsupervised text analysis. However, comprehensive evaluations of a topic model remain challenging. Existing evaluation methods are either less comparable across different models (e.g., perplexity) or focus on only one specific aspect of a model (e.g., topic quality or document representation quality) at a time, which is insufficient to reflect the overall model performance. In this paper, we propose WALM (Words Agreement with Language Model), a new evaluation method for topic modeling that comprehensively considers the semantic quality of document representations and topics in a joint manner, leveraging the power of large language models (LLMs). With extensive experiments involving different types of topic models, WALM is shown to align with human judgment and can serve as a complementary evaluation method to the existing ones, bringing a new perspective to topic modeling. Our software package will be available at https://github.com/Xiaohao-Yang/Topic_Model_Evaluation, which can be integrated with many widely used topic models.
@@ 1 Introduction
Although topic models have been time-tested for two decades, as an unsupervised technique, comprehensive evaluations of a topic model remain challenging (Zhao et al., 2021). Originally, topic models are implemented as probabilistic graphical models such as Latent Dirichlet Allocation (LDA) (Blei et al., 2003) and many of its Bayesian extensions. For these models, it has been common practice to measure the log-likelihood or perplexity of a model on held-out test documents. While log-likelihood or perplexity provides a straightforward quantitative comparison between models, several issues still persist: As topic models are not primarily designed for predicting words in documents but rather for semantically learning interpretable topics and document representations, these metrics do not measure these qualities. Furthermore, estimating the predictive probability is often intractable for Bayesian models, and different papers may employ different sampling or approximation techniques (Wallach et al., 2009; Buntine, 2009). For recently proposed neural topic models (NTMs) (Zhao et al., 2021), the computation of log-likelihood is even more inconsistent.
@@
In addition to log-likelihood or perplexity, document representation quality and topic quality are evaluated separately. For document representation quality, downstream task performance is typically used as a metric, such as document classification (Yang et al., 2023), clustering (Zhao et al., 2021), and retrieval (Larochelle and Lauly, 2012). For topic quality, the ultimate evaluation method is human evaluation, which is time-consuming and expensive. Thus, various automatic metrics have been proposed, such as topic coherence (Lau et al., 2014), which measures how semantically coherent the representative words in a topic are, and topic diversity (Dieng et al., 2020), which measures how diverse discovered topics are. To comprehensively evaluate the performance of a topic model, one needs to report multiple metrics on both document representation and topic qualities. However, these metrics can be contradictory; e.g., a topic model with good topic quality may not preserve good quality on document representation, and vice versa. This discrepancy complicates the model selection process for topic models in practice.
@@
In this paper, we aim to develop a new evaluation approach for topic modeling that comprehensively considers the semantic quality of document representations and topics in a joint manner, leveraging the power of large language models (LLMs). Our key idea is as follows: After being trained, a topic model can infer a document’s distribution over topics and each topic is a distribution over vocabulary words. With these two distributions, a model can generate a set of “topical” words given a document, such as by looking at its representative topics and the representative words of each topic. The generation of the topical words takes both the topic distribution of a document and the word distributions of the topics into account, which captures the semantic summary of the document and is expected to align with the keywords identified by humans. Thus, the quality of the generated topical words of documents informatively and comprehensively reflects the quality of a topic model. As human evaluation is typically costly, we propose querying LLMs with appropriate prompts as a proxy to generate keywords for the document and then compare them with the topical words generated by a topic model. Finally, to quantify the agreement between the words from the topic model and the LLM, a series of WALM (Words Agreement with Language Model) metrics are proposed. WALM has the following appealing properties:
@@
It is a joint metric that evaluates both document representation and topic qualities.
@@
It measures how well a topic model captures a document’s semantics, which is the focus of topic modeling.
It can be compared across different kinds of topic models.
@@ 2 Related Work
As for the evaluation of topics, Chang et al. (2009) design the word and topic intrusion tasks for human annotators, where high-quality topics or document representations are those where annotators can easily identify the intruders. Newman et al. (2010) and Mimno et al. (2011) evaluate topic coherence by direct ratings from human experts. Although human judgment is commonly regarded as the gold standard, it is expensive and impractical for large-scale evaluation. Automated evaluation of topic coherence is more practical, such as Normalized Pointwise Mutual Information (NPMI) (Lau et al., 2014), which relies on the co-occurrence of the topic’s top words in the reference corpus to measure topic coherence, with the underlying assumption that a large reference corpus such as Wikipedia can capture prevalent language patterns. Although they automate the evaluation of topics and strongly correlate with human judgment (Newman et al., 2010), counting word co-occurrence in a large reference corpus is still relatively expensive. Recent works propose leveraging word embeddings (Nikolenko, 2016) or contextualized embeddings (Hoover et al., 2021) for efficiently evaluating topic coherence, incorporating semantics from pre-trained embeddings. Due to common posterior collapse issues (Lucas et al., 2019) in the growing field of neural topic models (Zhao et al., 2021), recent works also consider topic diversity (Dieng et al., 2020) during evaluation, which measures how distinct the top words of each topic are.
@@
As for the evaluation of document representation, early works focus on how well the topic proportion of a document represents the document content, assessed through a topic intrusion task by human annotators (Chang et al., 2009), which is further extended as automated metrics (Bhatia et al., 2017, 2018). Recent topic models often use the topic proportions as document representations, the quality of which is commonly investigated through downstream tasks, including their use as features for document classification (Nguyen and Luu, 2021), clustering (Zhao et al., 2020), and retrieval (Larochelle and Lauly, 2012). Recently, the generalization ability of topic models is investigated by evaluating their quality of document representations across different unseen corpora (Yang et al., 2023).
@@
In the era of large language models (LLMs) (Thoppilan et al., 2022; Chowdhery et al., 2022; OpenAI, 2022; Touvron et al., 2023a, b), recent research has begun leveraging LLMs to evaluate topic models, such as using ChatGPT (OpenAI, 2022) as a proxy for human annotators for word intrusion and topic rating tasks for evaluating topic coherence (Rahimi et al., 2023; Stammbach et al., 2023). The focus of these works is still on topic quality only.
@@
In this work, we propose new evaluation metrics for topic models, differing from previous works in the following ways: (1) Unlike evaluations that focus on only sub-components of a topic model (i.e., topics or document representations), our evaluation metrics offer a joint approach to topic model evaluation, considering both topics and document representations together. (2) Compared with log-likelihood or perplexity, which also evaluate based on documents, our evaluation metrics consider semantics from documents and align with the focus of topic modeling. (3) Different from recent LLM-based evaluations that use LLMs for topic quality evaluation, ours is a more comprehensive method that considers both topic quality and document representation quality and our use of LLMs is quite different from previous works.
@@ 4.1 Motivation
Both topics and document representations are important components of a topic model. To comprehensively evaluate a topic model, it is common practice to report the performance of both parts. This can be done by measuring topic quality using metrics such as NPMI and assessing document representation quality through downstream classification accuracy (ACC) (see section 5.1 for details of metrics calculation). However, a model that prioritizes topic quality (e.g., NPMI) may not perform well in terms of document representations (e.g., ACC), and vice versa, which creates difficulty during model selection, as illustrated in Figure 1. This inconsistency in the performance of the two components is also indicated by Bhatia et al. (2017). Therefore, evaluating a topic model based on sub-components only is insufficient to reveal the entire model’s performance. Recent topic models usually focus on improving topic quality, such as clustering-based topic models in Sia et al. (2020); Grootendorst (2022), but they do not evaluate their performance for document representation. In this work, we aim to introduce a novel evaluation method for topic modeling that jointly assesses the semantic quality of both topics and document representations, with the help of large language models.
@@ 4.2 Key Idea
The inference process infers the document representation for a given document , as in Eq. 1; the generative process generates or reconstructs the word distribution based on the document representation and topics , as in Eq. 2. Therefore, the evaluation based on involves both document representations and topics.
Next, we take the top-weighted words from the word distribution generated by the topic model as the “topical” words of the document. Those topical words can be regarded as a semantic summary of the document from the target topic model’s perspective. To generate high-quality topical words for a document, a topic model should learn good global topics as well as good document representations. Now, the evaluation of the topic model can be converted to the evaluation of the quality of the topical words. Suppose the true representative words of document are given, then we can formulate our evaluation task as:
@@
4.3 Document summarization by LLM
Following our evaluation task in Eq. 4, the ideal representative words are from human summary of the document. However, this is expensive and impractical for large-scale evaluation. With the recent advancements in LLMs, which have demonstrated performance akin to human capabilities in various NLP tasks, including text summarization Wang et al. (2023); Tang et al. (2023); Zhang et al. (2024), we propose to leverage LLMs by prompting to extract key words of documents:
Specifically, we query keywords for a given document from an LLM by proposing the prompt shown in Figure 2. The prompt consists of a task instruction, an in-context example (Liu et al., 2021), and the queried document.
@@ 4.4 Choices of the Score Function
Considering the case that different words may describe the same or similar concept (e.g., both “puppy” and “dog”), we leverage the synset of WordNet (Miller, 1995) when counting the number of overlaps: if there is an overlap between two words’ synsets, then it is considered as an overlap, formulated as:
where denotes the indicator function; is a function that returns the synset for a given word.
@@
| subject to: | ||||
By finding the optimal binary matrix , we obtain the distance between and by:
@@
Considering that both and are top words of probability distributions, where each word essentially retains a portion of probability mass. Our previous calculations ignore the probability mass of words and treat each word in the set as equal. Now we include the probability mass and formulate the similarity calculation between and as an optimal transport (OT) problem: Given two discrete distributions and , where and are the supports of those two distributions; and are their related probability vectors222We assume is a uniform distribution over the keywords from the LLM. Thus, is a uniform probability vector.; following the same construction of cost matrix in the previous OA problem, the OT problem between and is defined as:
| subject to: | ||||
where is the transport plan, whose entry indicates the amount of probability mass moving from to . Similarly, by finding the optimal transport plan using solvers such as those in Flamary et al. (2021), the OT distance between and is obtained by:
@@ 5.1 Experimental Setup
We conduct experiments on 7 popular topic models from traditional probabilistic to recent neural topic models. (1) Latent Dirichlet Allocation (LDA) (Blei et al., 2003), the most popular probabilistic topic model that assumes a document is generated by a mixture of topics. (2) LDA with Products of Experts (PLDA) (Srivastava and Sutton, 2017), an early NTM that applies the product of experts instead of the mixture of multinomials in LDA. (3) Neural Variational Document Model (NVDM) (Miao et al., 2017), a pioneer NTM that uses a Gaussian as the prior distribution of topic proportions of documents. (4) Embedded Topic Model (ETM) (Dieng et al., 2020), an NTM that involves word and topic embeddings in the generative process. (5) Neural Topic Model with Covariates, Supervision, and Sparsity (SCHOLAR) (Card et al., 2017), an NTM that applies a logistic normal prior on topic proportions and leverages extra information from metadata. (6) Neural Sinkhorn Topic Model (NSTM) (Zhao et al., 2020), a recent NTM based on an optimal transport framework. (7) Contrastive Learning Neural Topic Model (CLNTM) (Nguyen and Luu, 2021), a recent NTM that uses contrastive learning to regularize document representations. We keep all these models’ default settings as suggested in their implementations. All experiments are conducted 5 times with different random seeds; mean and std values are reported.
@@
We evaluate topic models using WALM series metrics we propose in Sec. 4.4, including , , , and . As for the settings of WALM, we use the word embeddings from GloVe (Pennington et al., 2014) pre-trained on Wikipedia333https://nlp.stanford.edu/projects/glove/. We set the number of top words for a document generated by the topic model to 10; The LLM we use is LLAMA2-13b-chat (Touvron et al., 2023b), which has 13 billion parameters and is fine-tuned for chat completion; We use a 4-bit quantization444https://huggingface.co/TheBloke/Llama-2-13B-chat-GGML version of this model for memory efficiency and leverage the llama.cpp555https://github.com/abetlen/llama-cpp-python library for easy inference. Notably, our evaluation metrics can be used with any LLMs. We use LLAMA2 in our study because it is open-sourced and cost-efficient.
@@
We also evaluate topic models with existing commonly used metrics to compare with ours. (1) Topic Coherence and Diversity: we apply NPMI to evaluate topic coherence using Wikipedia as the reference corpus, done by the Palmetto package666https://github.com/dice-group/Palmetto (Röder et al., 2015). Following standard protocol, we consider the top 10 words of each topic and obtain the average NPMI score of topics by selecting the top 50% coherent topics. As for topic diversity (TD), we compute the percentage of unique words in the top 25 words of all topics, as defined in Dieng et al. (2020). (2) Document Representation Quality: we conduct document classification and clustering to evaluate the representation capability of topic models. As for classification, we train a Random Forest classifier based on the training documents’ representation and test the accuracy (ACC) in the testing documents, as in previous works such as Nguyen and Luu (2021). As for clustering, we conduct K-Means clustering based on test documents’ representation and report the Purity (KM-Purity) and Normalized Mutual Information (KM-NMI), as in previous works such as Zhao et al. (2020). (3) Perplexity: we use document completion perplexity (Wallach et al., 2009) to evaluate the predictive ability of topic models. We split each test document into two equal-length folds randomly. Then we compute the document completion perplexity (DC-PPL) on the second fold of documents based on the topic proportion inferred from the first fold, as in previous works such as Dieng et al. (2020).
@@
| LDA_C | drive disk scsi hard card controller mb floppy ide sale | |
@@
| LLAMA2 | driver format magneto-optical disc operating system communication incompatibility | |
| Human | driver disc drive computer hardware software memory formatting format incompatible | |
@@
| LDA_C | film played directed baseball league australian major drama football award | |
@@
| LLAMA2 | wrong world film australian directed ian pringle victoria melbourne | |
| Human | film movie directed directer australian melbourne victoria |

### v2

@@ Abstract
Topic modeling has been a widely used tool for unsupervised text analysis. However, comprehensive evaluations of a topic model remain challenging. Existing evaluation methods are either less comparable across different models (e.g., perplexity) or focus on only one specific aspect of a model (e.g., topic quality or document representation quality) at a time, which is insufficient to reflect the overall model performance. In this paper, we propose WALM (Word Agreement with Language Model), a new evaluation method for topic modeling that considers the semantic quality of document representations and topics in a joint manner, leveraging the power of Large Language Models (LLMs). With extensive experiments involving different types of topic models, WALM is shown to align with human judgment and can serve as a complementary evaluation method to the existing ones, bringing a new perspective to topic modeling. Our software package is available at https://github.com/Xiaohao-Yang/Topic_Model_Evaluation.
@@ 1 Introduction
Although topic models have been time-tested for two decades, as an unsupervised technique, comprehensive evaluations of a topic model remain challenging (Zhao et al., 2021a). Originally, topic models are implemented as probabilistic graphical models such as Latent Dirichlet Allocation (LDA) (Blei et al., 2003) and many of its Bayesian extensions, e.g., in Blei et al. (2010); Paisley et al. (2015); Gan et al. (2015); Zhou et al. (2016); Zhao et al. (2018b, a). For these models, it has been common practice to measure the log-likelihood or perplexity of a model on held-out test documents. While log-likelihood or perplexity provides a straightforward quantitative comparison between models, several issues still persist. Since topic models are not primarily designed to predict words in documents but rather to learn semantically meaningful topics and interpretable document representations, these metrics fail to capture these aspects. Furthermore, estimating the predictive probability is often intractable for Bayesian models, and different papers may employ different sampling or approximation techniques (Wallach et al., 2009; Buntine, 2009). For recently proposed Neural Topic Models (NTMs) (Zhao et al., 2021a), the computation of log-likelihood is even more inconsistent.
@@
In addition to log-likelihood or perplexity, document representation quality and topic quality are evaluated separately. For document representation quality, downstream task performance is typically used as a metric, such as document classification (Yang et al., 2023), clustering (Zhao et al., 2021a), and retrieval (Larochelle and Lauly, 2012). For topic quality, the ultimate evaluation method is human evaluation, which is time-consuming and expensive. Thus, various automatic metrics have been proposed, such as topic coherence (Lau et al., 2014), which measures how semantically coherent the representative words in a topic are, and topic diversity (Dieng et al., 2020), which measures how diverse discovered topics are. To comprehensively evaluate the performance of a topic model, one needs to report multiple metrics on both document representation and topic qualities. However, these metrics can be contradictory, e.g., a topic model with good topic quality may not preserve good quality on document representation, and vice versa. This discrepancy complicates the model selection process for topic models in practice.
@@
In this paper, we aim to develop a new evaluation approach for topic modeling that considers both the semantic quality of document representations and topics in a joint manner, leveraging the power of Large Language Models (LLMs). Our key idea is as follows: After being trained, a topic model can infer a document’s distribution over topics and each topic is a distribution over vocabulary words. With these two distributions, a model can generate a set of “topical” words given a document, such as by looking at its representative topics and the representative words of each topic. The generation of the topical words takes both the topic distribution of a document and the word distributions of the topics into account, which captures the semantic summary of the document and is expected to align with the keywords identified by humans. Given the high cost of human evaluation, we propose using LLMs as a proxy by employing appropriate prompts to generate keywords for the document, which are then compared with the topical words produced by a topic model. Finally, to quantify the agreement between the words from the topic model and the LLM, a series of WALM (Word Agreement with Language Model) metrics are proposed. WALM has the following appealing properties:
@@
It is a joint metric that evaluates the quality of both document representations and topics.
@@
It assesses how effectively a topic model captures the semantics of a document, which is a core objective of topic modeling.
It allows for comparisons across various types of topic models.
@@ 2 Related Work
As for the evaluation of topics, Chang et al. (2009) design the word and topic intrusion tasks for human annotators, where high-quality topics or document representations are those where annotators can easily identify the intruders. Newman et al. (2010) and Mimno et al. (2011) evaluate topic coherence by direct ratings from human experts. Although human judgment is commonly regarded as the gold standard, it is expensive and impractical for large-scale evaluation. Automated evaluation of topic coherence is more practical, such as Normalized Pointwise Mutual Information (NPMI) (Lau et al., 2014), which relies on the co-occurrence of the topic’s top words in the reference corpus to measure topic coherence, with the underlying assumption that a large reference corpus such as Wikipedia can capture prevalent language patterns. Although they automate the evaluation of topics and strongly correlate with human judgment (Newman et al., 2010), counting word co-occurrence in a large reference corpus is still relatively expensive. Moreover, coherence metrics can vary depending on the reference corpus, and there is no single “right” reference corpus that is suitable for all datasets Doogan and Buntine (2021). Recent works propose leveraging word embeddings (Nikolenko, 2016) or contextualized embeddings (Hoover et al., 2021) for efficiently evaluating topic coherence, incorporating semantics from pre-trained embeddings. Due to common posterior collapse issues (Lucas et al., 2019) in the growing field of neural topic models (Zhao et al., 2021a), recent works also consider topic diversity (Dieng et al., 2020) during evaluation, which measures how distinct the top words of each topic are.
@@
As for the evaluation of document representation, early works focus on how well the topic proportion of a document represents the document content, assessed through a topic intrusion task by human annotators (Chang et al., 2009), which is further extended as automated metrics (Bhatia et al., 2017, 2018). Recent topic models often use the topic proportions as document representations, the quality of which is commonly investigated through downstream tasks, including their use as features for document classification (Nguyen and Luu, 2021), clustering (Zhao et al., 2021b), and retrieval (Larochelle and Lauly, 2012). Recently, the generalization ability of topic models is investigated by evaluating their quality of document representations across different unseen corpora (Yang et al., 2023).
@@
In the era of Large Language Models (LLMs) (Brown et al., 2020; Thoppilan et al., 2022; Touvron et al., 2023a, b; Chowdhery et al., 2024), recent research has begun leveraging LLMs to evaluate topic models, such as using ChatGPT 111https://openai.com/index/chatgpt/ as a proxy for human annotators for word intrusion and topic rating tasks for evaluating topic coherence (Stammbach et al., 2023; Rahimi et al., 2024). The focus of these works is still on topic quality only.
@@
In this work, we propose new evaluation metrics for topic models, differing from previous works in the following ways: (1) Unlike evaluations that focus on only sub-components of a topic model (i.e., topics or document representations), our evaluation metrics offer a joint approach to topic model evaluation, considering both topics and document representations together. (2) Compared with log-likelihood or perplexity, which also evaluate based on documents, our evaluation metrics consider semantics from documents and align with the focus of topic modeling. (3) Different from recent LLM-based evaluations that use LLMs for topic quality evaluation, ours considers both topic quality and document representation quality and our use of LLMs is quite different from previous works.
@@ 4.1 Motivation
Both topics and document representations are important components of a topic model. To comprehensively evaluate a topic model, it is common practice to report the performance of both parts. This can be done by measuring topic quality using metrics such as NPMI and assessing document representation quality through downstream classification accuracy (ACC) (see section 5.1 for details of metrics calculation). However, a model that prioritizes topic quality (e.g., NPMI) may not perform well in terms of document representations (e.g., ACC), and vice versa, which creates difficulty during model selection, as illustrated in Figure 1. This inconsistency in the performance of the two components is also indicated by Bhatia et al. (2017). Therefore, evaluating a topic model based on sub-components only is insufficient to reveal the entire model’s performance. Recent topic models often focus on improving topic quality, such as clustering-based models (Sia et al., 2020; Grootendorst, 2022), but they do not evaluate their effectiveness in representing documents. In this work, we aim to introduce a novel evaluation method for topic modeling that jointly assesses the semantic quality of both topics and document representations, with the help of large language models.
@@ 4.2 Key Idea
The inference process infers the document representation for a given document , as in Eq. 1; the generative process222We omit topics in Eq. 3 as they are considered part of the parameters of the generative process . generates or reconstructs the word distribution based on the document representation and topics , as in Eq. 2. Therefore, the evaluation based on involves both document representations and topics.
Next, we take the top-weighted words from the word distribution generated by the topic model as the “topical” words of the document. Those topical words can be regarded as a semantic summary of the document from the target topic model’s perspective. To generate high-quality topical words for a document, a topic model should learn good global topics as well as good document representations. Now, the evaluation of a topic model can be reframed as assessing the quality of its topical words. Suppose the true representative words of document are given, then we can formulate our evaluation task as:
@@
4.3 Word Suggestion by LLM
Keywords Suggestion
Following our evaluation task in Eq. 4, the ideal representative words are from human summary of the document. However, this is expensive and impractical for large-scale evaluation. With the recent advancements in LLMs, which have demonstrated performance akin to human capabilities in various natural language processing tasks, including text summarization Wang et al. (2023); Tang et al. (2023); Zhang et al. (2024) and keyphrase extraction Song et al. (2023); Maragheh et al. (2023); Bai et al. (2024), we propose leveraging LLMs through prompting to generate keyword suggestions for a given document:
Specifically, we query keywords for a given document from an LLM by proposing the prompt shown in Figure 2. The prompt consists of a task instruction and the queried document.
Topic-Aware Keywords Suggestion
Analogous to the generation of topical words in topic modeling – where global topics of the document collections are identified first, followed by associated keywords for each given document – we propose prompting the LLM in a similar manner, ensuring it considers collection-level topics when providing keywords suggestion for each document, written as:
where denotes the set of topics of the text corpus. To obtain the collection-level topics for the corpus by the LLM, we follow the topic generation approach by Pham et al. (2024). Briefly, it leverages an LLM to iteratively identify new topics from each document. A subsequent refinement process then merges similar topics and removes those with low frequency. For further details on the topic generation process, we refer readers to Pham et al. (2024) (section 3.1).
Using these corpus-level topics, we prompt the LLM to generate keywords for each document in a two-stage process, considering the overarching themes of the collection. In the first stage, the LLM selects relevant topics for the target document from the corpus-level topics. In the second stage, we prompt the LLM to generate indexing words for the document based on each selected topic. The final set of keywords is obtained by merging the words generated for each selected topic. An example prompt and output for topic-aware keywords suggestion is shown in Figure 3.
@@ 4.4 Choices of the Score Function
Considering the case that different words may describe the same or similar concept (e.g., “puppy” and “dog”), we leverage WordNet (Miller, 1995) synsets to determine word overlaps: if the synsets of two words intersect, they are considered to overlap. Then, we define the synset overlap score as:
where denotes the indicator function; is a function that returns the synset for a given word. Intuitively, the synset-based score builds on the idea of word overlap, considering two words as overlapping if their synsets intersect, rather than requiring an exact match.
@@
subject to and . By finding the optimal binary matrix , we obtain the distance between and by:
@@
Optimal Transport (OT) has recently been used as a powerful geometric tool to measure the distance between distributions, with rich applications in machine learning and related areas Ge et al. (2021); Zhao et al. (2021c); Nguyen et al. (2021); Wang et al. (2022); Guo et al. (2022); Bui et al. (2022); Vuong et al. (2023); Zhao et al. (2023); Ye et al. (2024); Vo et al. (2024); Gao et al. (2024). Considering that both and are top words of probability distributions, where each word essentially retains a portion of probability mass. Our previous calculations ignore the probability mass of words and treat each word in the set as equal. Now we include the probability mass and formulate the similarity calculation between and as an Optimal Transport (OT) problem: Given two discrete distributions and , where and are the supports of those two distributions; and are their related probability vectors333We assume is a uniform distribution over the keywords from the LLM. Thus, is a uniform probability vector.; following the same construction of cost matrix in the previous OA problem, the OT problem between and is defined as:
subject to and ; is the transport plan, whose entry indicates the amount of probability mass moving from to . Similarly, by finding the optimal transport plan using solvers such as those in Flamary et al. (2021), the OT distance between and is obtained by:
@@ 5.1 Experimental Setup
We conduct experiments on 7 popular topic models from traditional probabilistic to recent neural topic models. (1) Latent Dirichlet Allocation (LDA) (Blei et al., 2003), the most popular probabilistic topic model that assumes a document is generated by a mixture of topics. (2) LDA with Products of Experts (PLDA) (Srivastava and Sutton, 2017), an early NTM that applies the product of experts instead of the mixture of multinomials in LDA. (3) Neural Variational Document Model (NVDM) (Miao et al., 2017), a pioneer NTM that uses a Gaussian as the prior distribution of topic proportions of documents. (4) Embedded Topic Model (ETM) (Dieng et al., 2020), an NTM that involves word and topic embeddings in the generative process. (5) Neural Topic Model with Covariates, Supervision, and Sparsity (SCHOLAR) (Card et al., 2018), an NTM that applies a logistic normal prior on topic proportions and leverages extra information from metadata. (6) Neural Sinkhorn Topic Model (NSTM) (Zhao et al., 2021b), a recent NTM based on an optimal transport framework. (7) Contrastive Learning Neural Topic Model (CLNTM) (Nguyen and Luu, 2021), a recent NTM that uses contrastive learning to regularize document representations. We keep all these models’ default settings as suggested in their implementations. All experiments are conducted 5 times with different model random seeds; mean and standard deviation values are reported.
@@
For the WALM settings, we use GloVe word embeddings pre-trained on Wikipedia (Pennington et al., 2014)444https://nlp.stanford.edu/projects/glove/ in our embedding-based metrics. For the LLM generation settings, we use LLAMA3-8B-Instruct555https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct for our main experiments. We employ greedy decoding during LLM generation to ensure deterministic outputs, setting the maximum number of generated tokens to 300. When prompting the LLM, we limit the number of generated keywords to 5. For topical words from the topic model, we select the top 10 weighted words from the document-word distribution for each given document.
@@
We also evaluate topic models with existing commonly used metrics to compare with ours. (1) Topic Coherence and Diversity: we apply NPMI to evaluate topic coherence using Wikipedia as the reference corpus, done by the Palmetto package666https://github.com/dice-group/Palmetto (Röder et al., 2015). Following standard protocol, we consider the top 10 words of each topic and obtain the average NPMI score of topics by selecting the top 50% coherent topics. As for Topic Diversity (TD), we compute the percentage of unique words in the top 25 words of all topics, as defined in Dieng et al. (2020). (2) Document Representation Quality: we conduct document classification and clustering to evaluate the representation capability of topic models. As for classification, we train a Random Forest classifier based on the training documents’ representation and test the accuracy (ACC) in the testing documents, as in previous works such as Nguyen and Luu (2021). As for clustering, we conduct K-Means clustering based on test documents’ representation and report the Purity (KM-Purity) and Normalized Mutual Information (KM-NMI), as in previous works such as Zhao et al. (2021b). (3) Perplexity: we use document completion perplexity (Wallach et al., 2009) to evaluate the predictive ability of topic models. We split each test document into two equal-length folds randomly. Then we compute the Document Completion Perplexity (DC-PPL) on the second fold of documents based on the topic proportion inferred from the first fold, as in previous works such as Dieng et al. (2020).
@@
LDA_C | drive, disk, scsi, hard, card, controller, mb, floppy, ide, sale | |
@@
LLM | formatting, magneto-optical, driver, disc, incompatible | |
LLM (Topic-Aware) | troubleshooting, formatting, incompatibility, magneto-optical, driver, disc, mounting | |
Human | driver, disc, computer, hardware, software, memory, formatting, incompatible | |
@@
LDA_C | film, played, directed, baseball, league, australian, major, drama, football, award | |
@@
LLM | world, film, australian, directed, victoria | |
LLM (Topic-Aware) | film, industry, production, cinema, entertainment | |
Human | film, movie, directed, director, australian, melbourne, victoria |